Solve each of the following problems. Be sure to show the equation used for the solution.
Geometry. The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by ptaylor(2198)

Let x=Width(W)
Then 2x+2=Length(L)
Perimeter(P)=2W+2L so our equation to solve is:
34=2x+2(2x+2) get rid of parens
34=2x+4x+4 subtract 4 from both sides
34-4=2x+4x+4-4 collect like terms
30=6x divide both sides by 6
x=5 in-------------------------------------width (W)
2x+2=2*5+2=12 in---------------------------length(L)
ck
34=2*5+2*12
34=34
Hope this helps----ptaylor